please help me find the Internal rate of return

New Heritage Private Label Total Production Division Data: Dolls Accessories Dolls Accessories Revenue ($ millions) Operating Income ($ millions) 80 14 26 5 $125 $7.5 4.4 0.5 2.3 0.3
initial Expenditures (S thousands) Upfront R&D Upfront Marketing Investment in Working Capital Property, Plant & Equipment Total 2010 $ 625 625 800 1.470 $3,520
Answer #1

Operating income = Total Revenue – Direct Costs – Indirect Costs

Internal rate of return is [(Profit - Cost) / cost] * 100

Profit = Operating expense = 7.5 $ million

Cost = 3520 thousands

Thus rate of return is = [(7500000 - 3520000) / 3520000 ] * 100 = 1.13*100 = 113.08%